OpenStudy (anonymous):

does this equation factor? y=-x^2+8x-14

OpenStudy (anonymous):

You can't normally factor this quadratic so you must use the quadratic formula. So for any given quadratic in the form:\[\bf ax^2+bx+c=0\]The roots "x", is given by:\[\bf x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] @soccergurl7991

OpenStudy (anonymous):

so it wont factor?

OpenStudy (goformit100):

for checking if it is can be factored.. has too simple steps. you first have to check that if ther is any type of manipulation could be done to get